consider the following set of inequalitis: y_>5-2.5x
y_<3-x
x,y_>0
the correct graphical representation of this set of inequalities is given by
1
Expert's answer
2021-05-31T06:55:52-0400

Solution:

Given,

"y\ge5-2.5x;
\\y\le3-x;
\\x,y\ge0"

Plot "y=5-2.5x;\ y=3-x" on graph.

Put "x=0,y=0" in given inequations, we get,

"y\ge5-2.5x\Rightarrow 0\ge5\Rightarrow False\Rightarrow \text{Shaded portion/area under this inequality does not include O(0,0).}
\\y\le3-x\Rightarrow 0\le3\Rightarrow True\Rightarrow \text{Shaded portion/area under this inequality includes O(0,0).}"

"x,y\ge0 \Rightarrow \text{Feasible region/area is in first quadrant only}"



This is required graphical representation of given inequalities and shaded region is the solution.